OpenStudy (loser66):

Show that for all a>0, \(lim_{n\rightarrow \infty}\dfrac {a^n}{n!}=0\) Please, help

OpenStudy (loser66):

@phi

OpenStudy (anonymous):

you can break it up... \[\frac{ a }{ n } \cdot \frac{ a^{n-1} }{ \left( n-1 \right)! }\]limit of first term is 0 as n -> infinity

ganeshie8 (ganeshie8):

You could also try any of your favorite convergence tests

OpenStudy (loser66):

This problem is proving one, not just find out the limit.

OpenStudy (loser66):

My attempt: Let \(X_n = \dfrac{a^n}{n!}\) \(\forall a>0\) \(X_n>0\) so that By archimedean theorem, \(X_n\) bounded below.

OpenStudy (anonymous):

limit of product is product of the limits

OpenStudy (loser66):

Which show lim of \(X_n\) exists Let try \(X_{n+1}\)

OpenStudy (anonymous):

how does it show that the limit exist? it just says it's bounded below. you would have to show that X_n is strictly decreasing.

OpenStudy (loser66):

\(X_{n+1}= \dfrac{a^{n+1}}{(n+1)!}=\dfrac{a}{n+1}*\dfrac{a^n}{n!}=\dfrac{a}{n+1}X_n\)

OpenStudy (anonymous):

and it is, after some k...

OpenStudy (loser66):

if n>a, then \(X_n\) decreases if n<a, then \(X_n\) increases

OpenStudy (loser66):

I think I have to prove both cases, not just find out the lim. Am I right?

OpenStudy (loser66):

Archimedean theorem shows lim X_n exist

OpenStudy (anonymous):

you have to break it up to get to the mototone decreasing part. for some n>k>a you'll have \[\frac{ a^{n-k} }{ n \cdot \left( n-1 \right) \cdots \left( n-k \right)}\cdot \frac{ a^k }{ k! }\]the left one is monotone decreasing because each factor, from \(\frac{a}{n-k}\) to \(\frac{a}{n}\) is less than 1 and it is bounded below by 0.

OpenStudy (anonymous):

likewise the term on the right is bounded below and above because it has a finite number of terms, all of which are greater than 0.

OpenStudy (anonymous):

i'm not familiar with the archimedian theorem... i'll check it out.

OpenStudy (loser66):

One more question: What is the logic on: if X_n is decreasing, and a = lim X_n, then \(lim_{n\rightarrow \infty} (X_{n+1}) = a\)

OpenStudy (anonymous):

because X_n has a limit and if it is a sequence, then the difference between successive terms must go to 0.

OpenStudy (anonymous):

that is, if X_n -> a => X_n - X_n-1 -> 0 as n -> infinity

OpenStudy (loser66):

Thanks for being patient to me. Please, make it clearer